# OM4OV: Leveraging Ontology Matching for Ontology Versioning

### v1

@@ Abstract.
Due to the dynamic nature of the semantic web, ontology version control is required to capture time-varying information, most importantly for widely-used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component for efficient ontology management, the growing size of ontologies and accumulating errors caused by manual labour overwhelm current OV approaches. In this paper, we propose yet another approach to performing OV using existing ontology matching (OM) techniques and systems. We introduce a unified OM4OV pipeline. From an OM perspective, we reconstruct a new task formulation, performance measurement, and dataset construction for OV tasks. Reusing the prior alignment(s) from OM, we also propose a cross-reference mechanism to effectively reduce the matching candidature and improve overall OV performance. We experimentally validate the OM4OV pipeline and its cross-reference mechanism using three datasets from the Alignment Evaluation Initiative (OAEI) and exploit insights on OM used for OV tasks.
PVLDB Reference Format:
PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at URL_TO_YOUR_ARTIFACTS.
1. Introduction
@@
Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications (Guarino et al., 2009). An ontology is not static, and the need for version control arises with its birth. While Web data is dynamic, any ontology that is used needs to undergo periodic revisions to keep up with growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation (Klein and Fensel, 2001). For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” (Haller and Polleres, 2020). This may cause undesirable deficiencies in downstream artefacts that conform to or reuse the ontology that is being changed, leading to severe non-compliance and incompatibility issues.
@@
The foundation of ontology versioning (OV) aims at distinguishing and recognising changes between different ontology versions. By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology, can apply the correct changes correspondingly (Klein and Fensel, 2001). While various methods for OV have been developed, one approach is to extend the ontology itself with internal version information. An ontology can be issued with a unique identifier (e.g. IRI) or a specific version number (e.g. owl:versionInfo) to be distinguished from other versions; each ontology entity may have a new triple to record its current status (e.g. owl:DeprecatedClass and owl:DeprecatedProperty); or every triple could be extended with a 4th dimension populated with a triple timestamp, similar to the RDF-star schema. Alternatively, version information can be recorded in change logs. Change logs can take the form of notebooks, an extensible markup language (XML), or a knowledge graph (KG). Nevertheless, maintaining version information in the ontology can be time-consuming and labour-intensive. Either extending the current schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or has a separate change log. In such cases, current approaches have limited capabilities in detecting the wrong version information or discovering missing version information.
@@
In this study, we investigate a lightweight and fully automatic version control approach for ontologies. We observe that the nature of OV is very similar to that of ontology matching (OM). Both are introduced to tackle interoperability between ontologies, with ontology entities (mainly classes and properties) serving as inputs. While OM is a well-studied problem (Otero-Cerdeira et al., 2015), a unified pipeline can help extend and reuse existing OM techniques and systems for OV tasks. However, these OM techniques and systems cannot be directly used in OV tasks. OM and OV have some important differences. (1) The input of OM is two distinct ontologies, whereas the input of OV is two different versions of one ontology. The output of OM is a set of entity mappings, whereas the output of OV is two sets: changed entities and unchanged entities. (2) OM concentrates on similarities between two entities, while OV addresses differences between versions of one entity. OM determines matched entities between two different ontologies, while OV determines and distinguishes add, delete, remain, and update entities between different versions in one ontology. (3) In addition, there is a dearth of lab-based datasets for OV evaluation. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none of them contain benchmarks designed for OV tasks.
@@
To address these challenges, our aim is to explore the complementary correlation between these two tasks and to develop a uniform approach to shifting OV tasks towards OM tasks. In this setting, these two tasks become interchangeable and share the same pipeline. Specifically, our key contributions include:
@@
  • We introduce a novel OM4OV pipeline. The pipeline leverages OM for OV with new task formulation, measurement, and a testbed that includes data for benchmarking.
@@
  • Drawing on OM practice, a novel cross-reference mechanism is proposed to optimise the OV candidate selection and improve the overall performance of OM.
@@
  • We implement the OM4OV pipeline as a proof-of-concept system and experimentally evaluate the system in the Ontology Alignment Evaluation Initiative (OAEI) OV testbed.
@@
The remainder of the paper is organised as follows. Section 2 reviews the literature on OV. Section 3 introduces our novel OM4OV pipeline, while Section 4 implements and evaluates the proposed pipeline. We discuss our implications and current limitations in Section 5 and Section 6. Section 7 concludes the paper.
2. Related Work
@@
One option is to include version information inside the ontology. The Simple HTML Ontology Extensions (SHOE) (Heflin and Hendler, 2000) uses the tag BACKWARD-COMPATIBLE-WITH to record version information. Klein et al. (2002) argue that a carefully-managed version numbering system embedded in the URI of the ontology (and therefore the fully expanded name of entities defined in the ontology) can minimise the impact of adopting updated versions because unchanged entities will be unaffected in practice. These approaches have largely been adopted by the later OWL Web Ontology Language (Bechhofer et al., 2004), where a set of annotation properties related to version information is defined. These include owl:versionInfo and owl:priorVersion to describe the version number of the ontology, owl:backwardCompatibleWith and owl:incompatibleWith to specify the entity’s compatible or incompatible corresponding entity in the previous version, and owl:DeprecatedClass and owl:DeprecatedProperty to declare the deprecated entities. Later, the ontology language $\tau$OWL (Zekri et al., 2016) was introduced to extend the OWL triple schema to quadruples to represent the versioning of concepts within an ontology. This idea is now incorporated in the new proposals for RDF 1.2, which allow time-varying information to be deduced from a temporal dimension within the quadruple (RDF-star Working Group, 2024).
@@
A second option is to create a separate version log to track version changes. Unlike traditional approaches that use an unstructured plain text file, the authors in (Plessers and De Troyer, 2005) propose a new approach that uses a version ontology with the change definition language (CDL) to create a version log. In (Cardoso et al., 2020), the authors construct an historical knowledge graph (HKG). Storing the version log in the knowledge graph not only avoids repetition, but also enables advanced search functions. The authors in (Sassi et al., 2016) argue that version logs may contain redundancy and inconsistent information. They propose a graph-of-relevance approach for interlinking different version logs and removing less relevant versions.
@@
While current approaches simply record human-generated version information, less attention has been paid to machine-generated version information. In other words, both previous approaches rely on the version information contained in or attached to the ontology. If such information is missing or incorrect, there is no way to automatically detect versioning of ontology concepts. In this study, we introduce a lightweight and fully automatic OV approach. Our approach advances by reusing existing OM systems and techniques for OV tasks, rather than creating a new OV framework from scratch. This paves the way for transferring well-studied OM solutions to OV tasks. With minor modifications required, we can reuse existing OM techniques and systems for OV tasks. To the best of our knowledge, our work is the first attempt to systemically analyse and utilise OM for OV.
3. OM4OV Pipeline
@@
Fig. 1 illustrates the overall OM4OV pipeline. Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), an OM task can be considered as finding an alignment ($A$) that contains a collection of mappings. Similarly, given an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), an OV task can be considered as finding an alignment (A) that contains a collection of mappings over the two different versions. However, while an alignment for OM only considers matched entities, OV tracks both matched and non-matched entities. Further, matched entities are composed of two subsets remain and update, and non-matched entities are composed of add and delete entities. In practice, we expect for OV that unchanged remain entities dominate.
Figure 1. OM4OV pipeline.
3.1. Task Formulation
@@
The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},c\geq s\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or another relation, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$ (Euzenat, 2007).
@@
(1) $A_{match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},c\geq s\}=A\odot\cup A\otimes$
@@
  1. (1)
    The remain and update entities are actually matched entities between different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions. If the naming convention changes, we will classify ontology entities that have changed names according to systematic convention as update entities, by modelling their confidence $c$ as a high number less than $s$, but close to it. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions. Therefore, $A\odot$ and $A\otimes$ are defined as:
    (2) $A\odot=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},c=1\}=A_{match}(c=1)$
    $A\otimes=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},c<1\}=A_{match}(c\neq 1)$
  2. (2)
@@
    (3) $A\oplus=\{e2\in O^{\prime}$ $\cap A_{non-match}\}=\{e2\in O^{\prime}\}\setminus A_{match}$
    $A\ominus=\{e1\in O$ $\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\}\setminus A_{match}$
3.2. Performance Measurement
Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision, recall, and F1 score. While precision measures matching correctness and recall measures matching completeness, there is an inherent trade-off between precision and recall. The F1 score offers a harmonic mean to balance the matching correctness and completeness. They are defined as:
(4) $Precison=\frac{|A\cap R|}{|A|}\quad Recall=\frac{|A\cap R|}{|R|}$
@@ 3.2. Performance Measurement
(5) $N(O)+N(O^{\prime})$ $=2\times(|A\odot|+|A\otimes|)+|A\oplus|+|A\ominus|$
@@
(6) $|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=-|\Delta A\oplus|=-|\Delta A\ominus|$
@@
    (7) $Recall\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}$
@@
    (8) $Precision\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$
@@
3.3. Dataset Construction
We propose an approach to constructing synthetic OV datasets from OM datasets. Figure 2 illustrates the generation of OM4OV datasets. Our approach is described in the following steps:
Figure 2. Generation of OM4OV datasets.
  1. (1)
@@ 3.3. Dataset Construction
  2. (2)
    There are four possible changes in OV tasks: remain, update, add, and delete. Each ontology entity in $O_{i}$ is randomly assigned one of these.
  3. (3)
@@
    For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$. Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$. By construction, we also have that the four sets are pairwise disjoint.
  4. (4)
@@
Unlike the original OAEI datasets for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks.
3.4. Pipeline Optimisation

### v2

@@ Abstract
Due to the dynamic nature of the semantic web, ontology version control is required to capture time-varying information, most importantly for widely-used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component for efficient ontology management, the growing size of ontologies and accumulating errors caused by manual labour overwhelm current OV approaches. In this paper, we propose yet another approach to performing OV using existing ontology matching (OM) techniques and systems. We introduce a unified OM4OV pipeline. From an OM perspective, we reconstruct a new task formulation, measurement, and testbed for OV tasks. Reusing the prior alignment(s) from OM, we propose a pipeline optimisation method called cross-reference (CR) mechanism to improve overall OV performance. We experimentally validate the OM4OV pipeline and the cross-reference mechanism in modified Ontology Alignment Evaluation Initiative (OAEI) datasets. We also discuss the insights on OM used for OV tasks, where some false mappings detected by OV systems are not actually false.
Index Terms:
ontology matching, ontology versioning
I Introduction
@@
Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications [1]. An ontology is not static, and the need for version control arises with its birth. While Web data is dynamic, any ontology that is used needs to undergo periodic revisions to keep up with growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation [2]. For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [3]. This may cause undesirable deficiencies in artefacts that conform to or reuse the ontology that is being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
@@
The foundation of ontology versioning (OV) aims at distinguishing and recognising changes between different ontology versions. By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology, can apply the correct changes correspondingly [2]. While various methods for OV have been developed, one approach is to extend the ontology itself with internal version information. An ontology can be issued with a unique identifier (e.g. IRI) or a specific version number (e.g. owl:versionInfo) to be distinguished from other versions; each ontology entity may have a new triple to record its current status (e.g. owl:DeprecatedClass and owl:DeprecatedProperty); or every triple could be extended with a 4th dimension populated with a triple timestamp, similar to the RDF-star schema. Alternatively, version information can be recorded in change logs. Change logs can take the form of notebooks, an extensible markup language (XML), or a knowledge graph (KG). Nevertheless, maintaining version information in the ontology can be time-consuming and labour-intensive. Either extending the current schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or has a separate change log. In such cases, current approaches have limited capabilities in detecting the wrong version information or discovering missing version information.
@@
In this study, we investigate a lightweight and fully automatic version control approach for ontologies. We observe that the nature of OV is very similar to that of ontology matching (OM). Both are introduced to tackle interoperability between ontologies, with ontology entities (mainly classes and properties) serving as inputs. While OM is a well-studied problem [4], a unified pipeline that is able to extend and reuse existing OM techniques and systems for OV tasks would be very desirable. However, OM techniques and systems cannot be directly used in OV tasks because OM and OV have some important differences. (1) The input of OM is two distinct ontologies, whereas the input of OV is two different versions of one single ontology. The output of OM is a set of entity mappings, whereas the output of OV is two sets: changed entities and unchanged entities. (2) OM concentrates on similarities between two entities, while OV addresses differences between versions of one entity. OM determines matched entities between two different ontologies, while OV determines and distinguishes add, delete, remain, and update entities between different versions in one ontology.
@@
To address these challenges, our aim is to explore the complementary correlation between these two tasks and to develop a uniform approach to shifting OV tasks towards OM tasks. Specifically, our key contributions include:
@@
  • We systematically analyse OV from an OM perspective and introduce a novel OM4OV pipeline. The pipeline leverages OM for OV with new task formulation, measurement, and a testbed that includes data for benchmarking.
@@
  • Drawing on OM practice, a novel cross-reference mechanism is proposed to optimise candidate selection and improve the overall performance of OV.
@@
  • We implement the OM4OV pipeline as a proof-of-concept system and experimentally evaluate its performance.
@@
The remainder of the paper is organised as follows. Section II reviews the literature on OV. Section III introduces our novel OM4OV pipeline, while Section IV implements and evaluates the proposed pipeline. We discuss our implications and current limitations in Section V and Section VI. Section VII concludes the paper.
II Related Work
@@
One option is to include version information inside the ontology. The Simple HTML Ontology Extensions (SHOE) [5] uses the tag BACKWARD-COMPATIBLE-WITH to record version information. The authors in [6] argue that a carefully-managed version numbering system embedded in the URI of the ontology (and therefore the fully expanded name of entities defined in the ontology) can minimise the impact of adopting updated versions because unchanged entities will be unaffected in practice. These approaches have largely been adopted by the later OWL Web Ontology Language [7], where a set of annotation properties related to version information is defined. These include owl:versionInfo and owl:priorVersion to describe the version number of the ontology, owl:backwardCompatibleWith and owl:incompatibleWith to specify the entity’s compatible or incompatible corresponding entity in the previous version, and owl:DeprecatedClass and owl:DeprecatedProperty to declare the deprecated entities. Later, the ontology language $\tau$OWL [8] was introduced to extend the OWL triple schema to quadruples to represent the versioning of concepts within an ontology. This idea is now incorporated in the new proposals for RDF 1.2, which allow time-varying information to be deduced from a temporal dimension within the quadruple [9].
@@
A second option is to create a separate version log to track version changes. Unlike traditional approaches that use an unstructured plain text file, the authors in [10] propose a new approach that uses a version ontology with the change definition language (CDL) to create a version log. In [11], the authors construct an historical knowledge graph (HKG). Storing the version log in the knowledge graph not only avoids repetition, but also enables advanced search functions. The authors in [12] argue that version logs may contain redundancy and inconsistent information. They propose a graph-of-relevance approach for interlinking different version logs and removing less relevant versions.
@@
While current approaches simply record human-generated version information, less attention has been paid to machine-generated version information. In other words, both previous approaches rely on the version information contained in or attached to the ontology. If such information is missing or incorrect, there is no way to automatically detect versioning of ontology concepts. In this study, we introduce a lightweight and fully automatic OV approach. Our approach advances by reusing existing OM systems and techniques for OV tasks, rather than creating a new OV framework from scratch. This paves the way for transferring well-studied OM solutions to OV tasks. With minor modifications required, we can reuse existing OM techniques and systems for OV tasks. To the best of our knowledge, our work is the first to systemically analyse and utilise OM for OV.
III OM4OV Pipeline
@@
Fig. 1 illustrates the overview of OM4OV pipeline. Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), an OM task can be considered as finding an alignment ($A$) that contains a collection of mappings. Similarly, given an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), an OV task can be considered as finding an alignment (A) that contains a collection of mappings over the two different versions. However, while an alignment for OM only considers matched entities, OV tracks both matched and non-matched entities. Further, matched entities are composed of two subsets remain and update, and non-matched entities are composed of add and delete entities. In practice, we expect for OV that unchanged remain entities dominate.
Figure 1: OM4OV pipeline.
III-A Task Formulation
@@
The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},c\geq s\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or other relations, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$ [13].
$\footnotesize A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},c\geq s\}$ (1)
@@
$A_{match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},c\geq s\}=A\odot\cup A\otimes$ (2)
@@
  1. 1.
    The remain and update entities are actually matched entities between different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions. If the naming convention changes, we will classify ontology entities that have changed names according to systematic convention as update entities, by modelling their confidence $c$ as a high number more than $s$ but less than 1. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions. Therefore, $A\odot$ and $A\otimes$ are defined as:
    $A\odot=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},c=1\}=A_{match}(c=1)$ (3)
    $A\otimes=\{(e1,e2,\equiv,c)|$ $e1\in O,e2\in O^{\prime},s\leq c<1\}=A_{match}(c\neq 1)$
  2. 2.
@@
    $A\oplus=\{e2\in O^{\prime}$ $\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$ (4)
    $A\ominus=\{e1\in O$ $\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
III-B Measurement
Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision, recall, and F1 score. While precision measures matching correctness and recall measures matching completeness, there is an inherent trade-off between precision and recall. The F1 score offers a harmonic mean to balance the correctness and completeness. They are defined as:
$Precison=\frac{|A\cap R|}{|A|}\quad Recall=\frac{|A\cap R|}{|R|}$ (5)
@@ III-B Measurement
$N(O)+N(O^{\prime})$ $=2\times(|A\odot|+|A\otimes|)+|A\oplus|+|A\ominus|$ (6)
@@
$|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=-|\Delta A\oplus|=-|\Delta A\ominus|$ (7)
@@
    $Recall\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}$ (8)
@@
    $Precision\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (9)
@@
III-C Testbed
There is a dearth of lab-based datasets for OV evaluation. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none of them contain benchmarks designed for OV tasks. We propose an approach to constructing synthetic OV datasets from OM datasets. Fig. 2 illustrates the generation of OM4OV datasets. Our approach is described in the following steps:
Figure 2: Generation of OM4OV datasets.
  1. 1.
@@ III-C Testbed
  2. 2.
    There are four possible changes in OV tasks: remain, update, add, and delete. Each ontology entity in $O_{i}$ is randomly assigned one of these. For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$. Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$. By construction, we also have the four sets as pairwise disjoints.
  3. 3.
@@
  4. 4.
@@
  6. 6.
    We also incorporate the cross-reference mechanism into our OV testbed, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules: (1) $R_{or}$ is the original reference.xml that removes all the mappings related to add entities. (2) $R_{o^{\prime}r}$ is the original reference.xml that removes all the mappings related to delete entities and updates all the mappings related to update entities.
Unlike the original OAEI datasets for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks. The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks. Table I lists the detailed information of the selected OAEI track used for the OV testbed. The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies.
TABLE I: Selected OAEI track for the OV testbed.
III-D Cross-Reference Mechanism